details of hybridisation of MnO4-

Grade:12th Pass

4 Answers

Ashwin Sinha
12 years ago

Dear Najam Uddin,

I = GA + 1/2(VE -V -C)

= 4 + 1/2(7 -4+1)

= 6

hybridisation = sp3d2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
Regards,
Askiitians Experts
Ashwin

Plz. approve my answer by clicking 'Yes' given below, if you loved it. Plz. Plz. Plz. don't forget to do it.Smile

pulipaka subbu
12 years ago
HYBRIDISATION  DEPICTING FORMULA IS
HYBRIDISATION=1/2{VALENCE ELECTRONS OF CENTRAL ATOM +TOTAL MONOVALENTLY BOND ATOMS-ANION CHARGE +CATION CHARGE}
THEREFORE FOR MnO4- H=1/2{7+0+1} = 4 {SP3}
Swadesh Vaibhav
7 years ago
MnO4- has d3s hybridization. Here, Mn is in +7 OS. The oxygen atoms act as ligands to the manganese center, and donate electrons to its 3 dxy, dyz and dzy orbitals and to 4s orbital, thus the hybridization is d3s. Feel free to ask more.
Shihani anthony
5 years ago
I think its d2sp hybridization.because this MnO4- has tetrahedral geometry and it has no lone pairs.only 3 double bond and a single bond.that means 4 sigma bonds and 3 pye bonds.so therefore 7 unpaired electrons should be there in the orbitals.in the excitaion one of two electrons of 4s is excited to the 4p orbital.then it gives 7 valance electrons to form bonds with 4 Oxygens.so we can get as hybrid orbitals 2 3d orbitals,4s orbital and 1 4p orbital for the hybridization.this is an idea got to think..if it is incorrect please let me know.thank you.